### v1

# Persistence Paradox in Dynamic Science

Abstract
Persistence is often regarded as a virtue in science. In this paper, however, we challenge this conventional view by highlighting its contextual nature, particularly how persistence can become a liability during periods of paradigm shift. We focus on the deep learning revolution catalyzed by AlexNet in 2012. Analyzing the 20-year career trajectories of over 5,000 scientists who were active in top machine learning venues during the preceding decade, we examine how their research focus and output evolved. We first uncover a dynamic period in which leading venues increasingly prioritized cutting-edge deep learning developments that displaced relatively traditional statistical learning methods. Scientists responded to these changes in markedly different ways. Those who were previously successful or affiliated with old teams adapted more slowly, experiencing what we term a rigidity penalty - a reluctance to embrace new directions leading to a decline in scientific impact, as measured by citation percentile rank. In contrast, scientists who pursued strategic adaptation - selectively pivoting toward emerging trends while preserving weak connections to prior expertise - reaped the greatest benefits. Taken together, our macro- and micro-level findings show that scientific breakthroughs act as mechanisms that reconfigure power structures within a field.
Keywords: strategic adaptation, rigidity penalty, status, innovation.
1 Introduction
Scientists are trained in specialized fields. While the ideal of the “Renaissance man” has long been celebrated in intellectual communities [22], the scientific system is increasingly marked by specialization, leading to a growing “burden of knowledge,” which makes it challenging for any single individual to fully grasp all aspects of their domain [29, 53, 59]. This change could drive the increasing reliance on collaboration in the scientific system [11, 48] as well as the decreasing disruption in contemporary scientific research [37]. Most scientists tend to focus on a limited number of research topics, despite recent trends toward exploration [64].
Against this backdrop, the debate between persistence and flexibility becomes central to understanding a researcher’s career trajectory. On the one hand, flexibility has long been considered useful for successful scientific careers. Science is often described as an “endless frontier” [17] requiring researchers to adapt to new questions, topics, methods, and data [26, 42, 51]. Collaboration, in particular, has proven to be a powerful mechanism for researchers to acquire the expertise needed to pivot their research directions [63, 54].
On the other hand, a contrasting view posits that researchers who concentrate intensively on a single niche and persistently build on their prior work are more likely to achieve long-term impact. Social systems, including scientific communities, are inherently resistant to change to some degree [16, 33]. Empirical evidence suggests that adaptive researchers or those without a clear disciplinary identity often face penalties in both the production and perception of their work [27, 35, 45, 55]. Furthermore, although these researchers usually achieve higher long-term impact, their works may receive delayed recognition [66, 35]. Recent works have uncovered the career-dependent nature of persistence, noting that frequent topic switching negatively affects early-career researchers’ productivity but positively impacts productivity later in their careers [64]. Slightly shifting one’s research focus can be beneficial for postdocs entering the faculty job market, even if this strategy is not universally adopted by all researchers [21].
In this paper, we contribute to the ongoing debate by offering new insights into how persistent researchers could lose impact under the paradigm shift of a field that is well represented in leading publishing venues. Our analysis takes place within a dynamic period — in particular, the transformative shift following the publication of AlexNet [1, 32], coincided with a time of rapid and unpredictable growth in the popularity of deep neural networks, which began to replace relatively traditional statistical machine learning methods  [^1] We will have more discussions about the history in the next section.. Our findings suggest that while persistence is often seen as a virtue in research, its value is highly context-dependent. In a rapidly evolving scholarly ecosystem, rigid adherence to an old niche can limit a researcher’s ability to remain relevant and influential. Conversely, strategic flexibility that balances adaptation and persistence enables researchers to align with emerging priorities, thereby enhancing their long-term career impact. These dynamics raise important questions about the trade-offs researchers face between cultivating depth and maintaining adaptability.
The remainder of this paper is organized as follows. In Section 2, we provide a concise historical overview of deep learning, emphasizing the period that forms the focal point of our study. Section 3 reviews related work and theoretical perspectives that underpin and motivate our analysis. In Section 4, we describe the datasets employed in our research. Section 5 investigates the adaptation mechanisms and their relationship to researchers’ success. We conclude with final remarks in Section 6.
2 A Brief History of Deep Learning
Over the past three decades, machine learning has undergone a profound transformation, marked by a paradigm shift from relatively traditional statistical methods to advanced deep learning models. These black-box models, characterized by their multi-layered architectures and vast numbers of parameters, have come to dominate the field due to their exceptional scalability and performance. This sweeping transition has been driven by the joint force of a large number of groundbreaking research and significant institutional changes, while with a widely acknowledged starting point – the seminal paper AlexNet [32], presented at the Neural Information Processing Systems conference in 2012. In 2015, three “founding fathers” of deep learning — Geoffrey Hinton, Yoshua Bengio, and Yann LeCun — authored a review paper titled “Deep Learning” published by Nature [36] and explicitly identified AlexNet as the catalyst for the start of the deep learning revolution. Since NeurIPS 2012 was held in December of that year, 2013 is regarded in our study as the first year marking the paradigm shift in the machine learning community. What makes this period particularly remarkable is not only the eventual success of deep learning, but the fact that the paradigm shift it triggered was largely unanticipated by the mainstream machine learning community even the authors themselves. Although foundational discussions around deep neural networks had taken place prior to AlexNet, the topic remained marginal, often dismissed as impractical due to its high computational demands (only later addressed through GPU acceleration) and perceived lack of theoretical elegance. Before AlexNet’s breakthrough performance on ImageNet in 2012 - becoming the first model to achieve a sub-25% error rate - many in the academic community were unfamiliar with its authors, let alone expecting the transformative impact that followed  [^2] https://www.pinecone.io/learn/series/image-search/imagenet/. This episode is now often celebrated as a powerful example of scientists staying true to their convictions (more details see “The History Began from AlexNet: A Comprehensive Survey on Deep Learning Approaches” [1]).
Our research investigates how the scientific community navigated this paradigm shift. It is unreasonable to expect that the entire scientific community immediately embraced deep learning following the publication of AlexNet. Communities closely tied to the field of deep learning would have faced immediate pressure to adapt or risk obsolescence, while those further removed may exhibit less pronounced responses. In 2013, the “year one” of deep learning, two pioneers in the field, Yann LeCun and Yoshua Bengio, along with others, established the International Conference on Learning Representations (ICLR), a conference dedicated to the emerging topic of deep learning, as the pioneers of deep learning felt the need to separate themselves from the broader machine learning community. Within a few years, ICLR grew to become one of the three most impactful conferences in machine learning, alongside the International Conference on Machine Learning (ICML) and Neural Information Processing Systems Conference (NeurIPS, known as NIPS before 2018), both of which were established in the 1980s. They are widely recognized as having the highest impact in the field  [^3] Accessed on Aug 27, 2024, from Google Scholar’s impact rank of AI venues. and mutually acknowledge each other’s prominence. For example, the ICLR conference guidelines state that potential reviewers must be qualified by having at least one accepted publication in a prior ICLR, NeurIPS, or ICML conference proceeding  [^4] https://iclr.cc/Conferences/2025/CallForPapers. Accessed on Dec 31, 2024. In computer science, papers in a conference proceeding are considered formal publications..
Among publishing venues in AI and machine learning, ICLR can reasonably be considered the conference “closest” to deep learning in terms of focus, or “purity,” given its explicit connection to the topic. However, we do not expect the new conference to have published the most influential deep learning papers during its early years. ICML and NeurIPS, as leading conferences predating the deep learning revolution, likely represent a second layer of proximity to deep learning research. Papers published by these two conferences laid a solid foundation for the explosion of deep learning. As top-tier conferences, they have consistently been attuned to advancements at the scientific frontier and have emphasized state-of-the-art research. Furthermore, many founding members of ICLR were key contributors to ICML and NeurIPS during the pre-deep-learning era, underscoring the close intellectual ties among these venues. Other academic forums, such as specialized conferences in general artificial intelligence, statistics, natural language processing, and computer vision, undoubtedly feature high-quality deep learning research. However, these venues are less directly connected to the core of deep learning and therefore constitute a more distant layer of proximity in our analysis. The empirical evidence of these distances can be found in Appendix A.
3 Institutional and Individual Responses
to New Paradigms
3.1 Institutional Responses to New Paradigms
The scientific enterprise oscillates between periods of stability and episodes of disruptive change. During periods of “normal science,” research communities adhere to shared paradigms and incrementally extend them, creating a stable hierarchy of knowledge and authority [33]. Over time, however, unresolved anomalies accumulate and radical new ideas – often initially developed at the margins – precipitate a crisis leading to a paradigm shift [7, 28, 10]. Such disruptive paradigm shifts can upend existing hierarchies and reward structures in science. Established leaders and institutions, heavily invested in the old paradigm, may resist the new one even as evidence mounts. In effect, the hierarchy is overturned as younger scholars embracing the new framework eventually replace the old guard. However, recent empirical research paints a more nuanced picture of how disruption unfolds. When a dominant figure in a field is removed or a prevailing theory collapses, it can create an opening for innovation by younger generations [7, 4]. In essence, the loss of a field’s apex leader loosened the paradigm’s hold, allowing diverse approaches to flourish until the field’s overall productivity recovered and even exceeded the previous level. This supports a variant of Planck’s principle [28]: rather than age alone, it is often the position and influence of incumbents that shape inertia. As long as a dominant elite sets research agendas (consciously or not), outsiders struggle to introduce disruptive innovations. When that hold is broken – whether by generational turnover, a scientific crisis, or external shock – hierarchical barriers give way to new ideas and players [13, 31, 46, 18].
Institutional inertia - including conferences, journals, universities, and funding bodies – plays a critical role in mediating the tension between persistence and adaptation during paradigm shifts. These institutions often embody the prevailing consensus of the current paradigm and therefore tend to respond conservatively to radically new research [34], while being more receptive to innovations that align closely with existing agenda [62, 61]. For instance, elite journals and conference program committees are typically staffed by established experts whose standards were forged under the prevailing paradigm. Through the norm of organized skepticism, they subject unconventional findings to especially stringent scrutiny [39, 34, 52]. This rigorous peer review serves as quality control, ensuring that only well-substantiated paradigm-challenging results enter the literature. However, it can also slow down the acceptance of novelties: truly revolutionary work might face higher hurdles to publication, as reviewers demand extensive evidence to be convinced of claims that contradict textbook knowledge [52, 5]. Reviewers even care more about how the work is positioned and connected with prior literature than its substantive contribution [9, 50]. In the early phase of a paradigm shift, researchers aiming to publish paradigm-breaking studies may seek out alternative outlets. We often observe the emergence of new journals, special issues, or conference tracks dedicated to the innovative approach, providing a collegial arena for like-minded scholars to develop the paradigm. Such specialized forums reduce the penalization of deviant ideas and allow the new paradigm to build credibility through the accumulation of results. Over time, if the paradigm gains adherents and empirical support, mainstream institutions adapt [10]. Prestigious journals may start accepting (and even soliciting) papers from the new paradigm [52, 6, 5, 24], especially as influential scientists convert or a new generation of editors takes the helm. Major conferences might incorporate sessions on the once-controversial topic, signaling its integration into the core disciplinary dialogue. This institutional evolution can be seen as a shift in the “logic” of the field’s reward system: what was once viewed as heretical gradually becomes high-impact cutting-edge research. Furthermore, academic funding agencies and professional societies adjust their priorities in response to paradigm changes. They might create grant programs targeting the new paradigm’s questions, or give awards and keynote invitations to its pioneers, thereby realigning incentives for individual researchers [3, 6]. Throughout this process, there is an interplay between persistence and flexibility: institutions must uphold rigorous standards and continuity (to avoid credulously endorsing every novelty), yet they also must remain open to creative destruction in science. The long-term vitality of a discipline arguably depends on this balance – allowing new paradigms to thrive when evidence warrants, while preserving enough stability for cumulative knowledge building. Thus, conferences and journals act as both gatekeepers and enablers: initially gatekeeping by applying skepticism to unorthodox ideas, but eventually becoming enablers of a new paradigm once its robustness is demonstrated and a critical mass of the community has shifted its consensus.
3.2 Individual Responses to New Paradigms
Individual scientists face an “essential tension” between exploiting established expertise and exploring novel ideas [57]. One can either refine known competencies or venture into uncertain territory for potential breakthroughs. Scientists are more likely to pursue incremental, “normal science” projects than to tackle novel questions that create new connections between ideas [44, 23]. This conservative tendency persists even as new opportunities expand, suggesting many scientists stick to narrow paradigms that reliably yield publications. Such persistence can bolster short-term productivity and career stability, but it may come at the cost of innovation. Crucially, the reward structures in science often reinforce conservative behavior. In principle, Mertonian norms of communalism, universalism, disinterestedness, and organized skepticism encourage openness to new ideas. However, the institutional reward system of science – funding, tenure, and prestige – tends to favor cumulative productivity and established reputations [56]. Once a researcher achieves recognition or funding, they become more likely to gain further success, creating a cycle of cumulative advantage. This dynamic can make established scientists reluctant to risk their status by deviating into new fields, while early-career researchers feel pressure to produce steady publications rather than high-risk breakthroughs. In other words, persistence in a proven line of research is often the rational career strategy in an environment where reviewers, funding agencies, and tenure committees reward incremental progress.
Despite these pressures toward specialization, some researchers do embrace flexibility and adaptation. Prize-winning and highly innovative scientists tend to introduce novel combinations of ideas far more often than their peers [60]. While most scientists repeatedly tested known relationships, a minority pursued risky new links between distant knowledge “nodes.” Those novel studies, when they succeeded, garnered higher citation impact on average (i.e. they were more influential). This high-risk, high-reward pattern suggests that adaptable scientists who recombine knowledge from diverse areas can drive disproportionately impactful discoveries – even though they face greater odds of failure or initial rejection. Indeed, Nobel laureates and other esteemed scientists have often been more adventurous, switching topics or methods to follow emerging questions. Such adaptability may slow one’s early career (as new skills or credibility must be built), but it can pay off in path-breaking contributions later [64]. In other words, the breadth of research can dilute short-term impact even as it expands long-term innovative capacity. This underscores the personal tension researchers navigate between deep persistence in one area and flexible exploration across domains.
Cognitive and social factors also determine why some scientists adapt while others persist. Cognitive lock-in and expertise traps can make a veteran researcher deeply invested in a particular paradigm less likely to “see” merit in new approaches [8]. By contrast, newcomers or those with interdisciplinary training might more readily apply fresh frameworks. Age is also a predictor of openness [4]. In some historical cases – for example, the acceptance of plate tectonics in geology – senior scientists actually adopted the new paradigm faster than their juniors [20, 40]. Such cases suggest that established experts can and do pivot under the right conditions, perhaps due to broader experience or fewer career risks at late stages. Other evidence shows that across all fields, time periods, and impact levels, scientists increasingly rely on older ideas and references as their careers progress [19]. Nonetheless, the costs of switching fields or approaches (in time required to retrain, loss of status in one’s old community, uncertainty of recognition in the new area) mean that many researchers only make bold shifts when compelled by clear evidence or lack of progress in their current line. Those who do successfully adapt often leverage transferrable skills to ease the transition – for instance, applying a known method to a new domain [2].
Who will be more adaptive? One line of theory argues that low-status scientists are more inclined to change their research directions during periods of scientific disruption. Limited access to resources, recognition, and opportunities within the existing paradigm creates structural constraints that may push low-status scientists to explore alternative paths [14, 15, 25, 49]. Disruptive events, which destabilize established hierarchies, may offer unique opportunities for these scientists to reposition themselves. By breaking from dominant paradigms, they can seek new avenues for development, effectively leveraging disruption to challenge the entrenched status quo [65], while successful scientists are less likely to adapt because they are reluctant to let go of their existing achievements. Sociologists have extensively examined the interplay between disruption and inequality. While disruptions may exacerbate existing inequities (e.g., by amplifying gaps), they also hold the potential to redistribute opportunities for those willing to adapt [65, 58]. Low-status scientists, in particular, may experience a heightened motivation to adjust their research agendas during these periods, as they face fewer risks associated with deviating from the dominant scientific order and they may gain status during the process [38].
Conversely, another perspective posits that high-status scientists, by virtue of their abundant resources and influence, are more likely to first adapt during periods of scientific paradigm shifts. These scientists are often positioned at the forefront of their fields, leveraging their significant access to research funding, cutting-edge experimental infrastructure, and extensive academic networks to adapt to new paradigms [7, 39]. ICLR illustrates how prominent researchers in machine learning mobilized their resources and networks to institutionalize and promote advancements in deep learning [36]. In addition to the prior status of scientists, the collaborator network is also likely to have a significant influence on an author’s research adaptation. Strong collaborators may lead to greater changes toward their research.
Research is a deeply social process: new ideas spread through formal and informal networks (colleagues, co-authors, mentors, conference communities), and a researcher’s status and position in these networks can either constrain or enable adaptation. Overall, at the micro-level, a scientist’s decision to persist or pivot emerges from the interaction between collaborations, personal curiosity, risk tolerance, career incentives, and the perceived credibility of the emerging paradigm.
4 Dataset
To examine how science adapts during a transformative era, we focus on researchers active before the deep learning boom in 2013. We begin by collecting all 22,651 papers published at NeurIPS and ICML in the ten years before and after 2013 (2003-2022) and 5,163 ICLR papers published during 2013-2022, capturing shifts in the machine learning landscape. We also assemble a cohort of 5,359 scholars who published at least one paper in ICML or NeurIPS between 2003 and 2012—i.e., the decade preceding the deep learning surge—and construct a panel dataset tracking their annual research output at those same conferences from 2003 through 2022 (ten years before and after the deep learning boom).
We did not use these authors’ papers back to the earliest stages of machine learning in the 1980s. Given the generational turnover of scientists and the evolutionary nature of scientific progress, their papers between 2003 and 2012 are more representative of the pre-deep learning scientific community and the immediate reaction of the community to a paradigm shift. These scientists were relatively senior when they encountered the deep learning boom in 2013. We recorded their academic age—calculated as the difference between 2013 and the year of their first publication in the OpenAlex database—to represent their experience at the time of this paradigm shift (Figure 1). On average, they had been active in the field for over a decade, having published in the most prestigious venues during the formative years of 2003–2012.
Figure 1: The seniority of scientists at the time of the machine learning paradigm shift.
Our central research setting is the group of machine learning researchers who published in top-tier conferences (ICML and NeurIPS) over the decade leading up to the advent of deep learning. When deep learning began its explosive growth in 2013, would these researchers - if they continued pursuing their established topics and methods - still be able to publish at the same premier venues, given the rapid evolution of the field, which is arguably well reflected in these conferences  [^5] We have more discussions in Appendix A.? Furthermore, how would their productivity and the impact of papers change compared to other papers appearing in ICML and NeurIPS during this transformative period?
The purpose of selecting this sample is two-fold: (a) ICML and NeurIPS are premier conferences with a plausibly high bar of quality and visibility. They are particularly relevant to the deep learning revolution and have the advantage of existing both before and after 2013, making them ideal as a reference for longitudinal analysis. (b) If a scientist lowers their standards for where they submit papers, their productivity will almost always increase, regardless of whether they actually adapt their work. In Appendix A, we analyze 579,973 papers from the Microsoft Academic Graph (MAG) dataset, categorized under ”machine learning” or ”deep learning,” spanning the same 20-year period of deep learning formation and growth. Using advanced embedding and computational linguistics techniques, we empirically demonstrate that, compared to ICML and NeurIPS, general machine learning publications are less closely tied to the rise of deep learning. In other words, these broader publications lack a consistent standard for “selecting” cutting edge work, or their submitting authors are less sensitive to the change, limiting their ability to provide a clear analysis of the evolution of fields and scientists.
We would like to highlight two key aspects of research design. First, name disambiguation is particularly crucial in computer science papers due to the high prevalence of common Chinese names [30] and the name disambiguation from OpenAlex is not perfect [67]. In our study, the scientists in our sample published at least one paper in ICML or NeurIPS between 2003 and 2012 - a period during which Chinese scientists were less represented in these conferences compared to their prevalence today. We use names, email addresses, and affiliations for initial name disambiguation. However, this approach is still limited, particularly in cases where individuals change their email addresses/institutions after graduating. To improve accuracy, we further utilize the Google Scholar API, Scholarly, for disambiguation. If two authors (with different email addresses or affiliations) have the same name, and their publications on Google Scholar indicate they are the same individual, we merge them into a single author record. Second, a significant confounder in our research design is the establishment of ICLR in 2013, which quickly rose to become a top-tier conference alongside ICML and NeurIPS. This development likely influenced the publishing intentions and patterns of our focus group in ICML and NeurIPS. However, in Appendix C, we empirically demonstrate that ICLR’s impact was relatively modest, at least during its initial years.
5 Adaptation and Success
This section presents our analysis of the relationship between adaptation and success within the evolving academic landscape. We begin with a macro-level perspective on how the field is changing (Section 5.1), followed by a micro-level analysis (Section 5.2), which explores whether adaptation contributes to success (Section 5.2.1) and who is more likely to adapt (Section 5.2.2). Finally, we briefly summarize the key insights from our results (Section 5.3).
5.1 Macro-level Analysis
How do ICML and NeurIPS capture the cutting edge of deep learning? ICLR was founded by the pioneers of deep learning to establish a dedicated venue for this rapidly growing field, making its research highly representative of deep learning. Using the TF-IDF (Term Frequency-Inverse Document Frequency) algorithm, we extracted the top ten keywords from each paper’s abstract. TF-IDF is a statistical method widely used to evaluate how important a word is to a document (abstract) within a collection of corpus (abstracts). The TF component measures how often a term appears in an abstract, capturing its relevance within that context. The IDF component, on the other hand, measures how unique or rare the term is across all abstracts in a year. By combining these two factors, TF-IDF assigns higher weights to terms that are frequent within one abstract but relatively rare across the entire corpus, allowing it to effectively highlight keywords that best represent the focal abstract’s unique topics. This approach provided a robust representation of the topics dominating ICLR each year and allowed for meaningful comparisons with topics emphasized at ICML and NeurIPS. An example of an abstract and its extracted tokens can be found in Appendix F.
Figure 2: Topical convergence between ICML, NeurIPS, and ICLR. Panel (a): the topical importance gap has been decreasing over the years. Panel (b): the ICLR-specific topics have been becoming more and more important in ICML and NeurIPS. Error bars represent 95% confidence intervals.
For every topic identified in ICLR in a given year, we calculated its frequency in ICML and NeurIPS abstracts from the same year (assigning a frequency of 0 if absent). We then computed the distribution of absolute differences in topic importance (frequency) between ICML and NeurIPS corpus versus ICLR corpus for that year. As shown in panel (a) of Figure 2, a clear trend emerges: the topics of the three conferences have become increasingly similar over time. Next, we identified the deep learning topics most underrepresented in ICML and NeurIPS, focusing on the top 50 topics with the largest differences in frequency between ICLR and ICML/NeurIPS during the first decade of the deep learning wave. These topics represent areas significantly more emphasized in ICLR than in the other two conferences. How have these topics evolved within ICML and NeurIPS? As shown in panel (b) of Figure 2, their importance began to rise following the onset of the deep learning wave. Moreover, the relative importance of these topics has converged among ICLR, NeurIPS, and ICML. These findings suggest that ICML and NeurIPS, either actively (through peer review and selection) or passively (through cross-conference submissions), have adapted to increasingly capture the cutting edge of deep learning.
5.2 Micro-level Analysis
5.2.1 Will more adaptive scientists be more successful?
In this section we study how adaptation contributes to success at the scientist level.
Dependent Variables (Academic Success):
  • Productivity: The number of papers published by a scientist in ICML and NeurIPS in a given year.
  • Impact (0-100): For each paper, we calculate its citation percentile rank among all papers published in the same conference that year. A scientist’s annual impact is the average percentile rank of all their papers for that year.
Independent Variables:
  • Research Persistence (0-100): To quantify a scientist’s research persistence from year $T$ to $T$ + 1, we first gather all ICML/NeurIPS papers authored by the scientist up to and including year $T$. Next, we collect their ICML/NeurIPS papers from year $T$ + 1. To assess persistence, we compute the bootstrapped text similarity between these two corpora. Since the sizes of the two corpora (before $T$ and $T$ + 1) may differ significantly—introducing potential bias (text similarity is typically lower when comparing a short text to a longer one)—we address this by matching their sizes. Specifically, we retain the smaller corpus and randomly sample a subset of texts from the larger corpus to match the size of the smaller corpus. We then calculate the text similarity between the sampled subset and the smaller corpus. This process is repeated 1,000 times, and the average similarity across iterations is used as the similarity score from year $T$ to $T$ + 1. A higher bootstrapped similarity score indicates greater persistence and less adaptation.
Control Variables:
We control the prior performance of scientists and the contributions of their collaborators. The core idea of our analysis is that a scientist’s success in a given year can be decomposed into three key components: (1) their individual ability to produce papers (measured by their prior performance), (2) the contributions of their collaborators (how many collaborators, how well the focal author knows them, and how good the collaborators are), and (3) the premium, or penalty, associated with research adaptation as determined by the publishing market, if any, which is our key independent variable. Adaptation could occur in part through collaboration. By controlling for components from collaboration networks, we isolate the additional impact that adaptation contributes.
  • Previous Productivity/Previous Impact (0-100): We use the author’s performance in the previous year to measure their prior foundation to publish.
The input from collaborators can be characterized by the number of collaborators, the strength of the focal author’s familiarity with them, the quality of those collaborators, and their potential interactions.
  • Collaborator Count: The total number of collaborators a scientist worked with on ICML/NeurIPS papers in a given year.
  • Collaboration Familiarity: For each scientist, we track their unique collaborators in a given year and calculate the number of prior ICML/NeurIPS papers they coauthored with these collaborators (dating back to the inception of these conferences: ICML 1980 and NeurIPS 1987). The scientist’s collaboration age for that year is defined as the average number of prior collaborations with their current collaborators.
  • Collaborator Status (0-100): Collaborator status is measured as the average status of all collaborators in year $T$. The status of a collaborator in year $T$ is determined by their percentile rank on the cumulative citations of produced ICML NeurIPS papers before year $T$ among all authors.
Descriptive statistics for the variables presented can be found in Figure 3. It reveals a long-tail distribution in productivity and collaboration networks. Interestingly, there is generally low persistence observed among authors in the field of computer science. The relationship between variables of persistence, prior status, age, and collaboration network features is shown in Appendix D.
Figure 3: Descriptive statistics for variables used in the regression.
Regression:
We track the cohort of scientists over 20 years using the regression with two-way fixed effects:
$Impact/Productivity_{i,t}=$
[…]
$+\alpha_{i}+\gamma_{t}+\varepsilon_{i,t}$
Where $\beta_{4}$$\beta_{6}$ terms represent the input from the collaborator network. $\varepsilon_{i,t}$ is the error term. Author fixed effects $\alpha_{i}$ are used to capture author-level characteristics, such as the scientist’s age, motivation, and learning ability when they face paradigm shifts. Year fixed effects $\gamma_{t}$ capture temporal trends and external factors that may influence publication outcomes in a given year, such as shifts in funding priorities, technological advancements, or broader disciplinary trends. These fixed effects help isolate variations due to the research environment from those attributable to individual scientists.
The model includes research persistence ($\beta_{1}$) and its interaction with the post-deep learning era ($\beta_{2}$) to examine how a scientist’s ability to adapt their research influences outcomes in the context of significant technological shifts. ${Research\ Persistence}_{i,t}$ captures the scientist $i$’s adaptation and persistence from the period before year $T$ to year $T$. These terms enable us to assess whether adaptivity offers additional benefits, or penalties, following the advent of deep learning.
In our regression results shown in Tables LABEL:impact_table and LABEL:prod_table (models 4, 5, and 6) we also tried models including interactions between features of the coauthor network in year $T$, Familiarity $\times$ Count, Familiarity $\times$ Status, Status $\times$ Count, and Status $\times$ Count $\times$ Familiarity, to control the potential synergistic effects of the familiarity of coauthors, coauthor status, and coauthor count on scientific outcomes. In all regressions standard error is clustered at the author level.
Table 1: Dependent variable: impact
Dependent variable: impact
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6
(persistence) (author) (collab. network)
post 2013$\times$×persistence -0.163∗∗∗ -0.204∗∗∗ -0.177∗∗∗ -0.177∗∗∗
(0.049) (0.051) (0.052) (0.052)
persistence 0.036 0.004 -0.002 -0.002
(0.045) (0.046) (0.046) (0.046)
prev. impact -$0∗∗$080∗∗∗ -$0∗∗$080∗∗∗
(0.014) (0.013) (0.013) (0.013)
collab. familiarity -4.172∗∗∗ -3.483∗∗∗ -2.028 -2.096
(0.590) (0.638) (1.034) (1.148)
collab. status 0.116∗∗∗ 0.116∗∗∗ 0.129∗∗∗ 0.128∗∗∗
(0.013) (0.015) (0.021) (0.022)
collab. count 0.297∗∗∗ 0.450∗∗∗ 0.711∗∗∗ 0.702∗∗∗
(0.063) (0.071) (0.111) (0.149)
familiarity×$\times$×status -0.006 -0.004
(0.017) (0.022)
familiarity×$\times$×count -0.393∗∗ -0.366
(0.121) (0.270)
status×$\times$×count -0.003 -0.003
(0.003) (0.004)
familiarity×$\times$×count×$\times$×status -0.001
two-way fixed effects Y Y Y Y Y Y
Observations 11,347 11,159 16,294 11,016 11,016 11,016
N. of groups 2,794 2,754 5,227 2,739 2,739 2,739
$R^{2}$ 0.004 0.005 0.017 0.029 0.031 0.031
Residual Std. Error 1.433 1.664 3.154 3.965 4.093 4.094
F Statistic 16.323∗∗∗ 43.563∗∗∗ 64.564∗∗∗ 41.696∗∗∗ 29.669∗∗∗ 26.701∗∗∗
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Table 2: Dependent variable: productivity
Dependent variable: productivity
Model 1 Model 2 Model 3 Model 4 Model 5 Model 6
(persistence) (author) (collab. network)
post 2013$\times$×persistence 0.037∗∗∗ 0.009∗∗∗ 0.005∗∗ 0.005∗∗
(0.004) (0.002) (0.002) (0.002)
persistence 0.031∗∗∗ 0.021∗∗∗ 0.023∗∗∗ 0.023∗∗∗
(0.002) (0.002) (0.002) (0.002)
prev. productivity 0.376∗∗∗ 0.034 -0.017 -0.010
(0.034) (0.021) (0.017) (0.017)
collab. familiarity 0.110∗∗∗ -0.028 -0.270∗∗∗ -0.528∗∗∗
(0.019) (0.021) (0.050) (0.073)
collab. status -0.001 -0.001∗∗ -0.001 -0.004∗∗
(0.000) (0.000) (0.001) (0.001)
collab. count 0.227∗∗∗ 0.197∗∗∗ 0.162∗∗∗ 0.126∗∗∗
(0.010) (0.009) (0.015) (0.017)
familiarity×$\times$×status -0.000 0.006∗∗∗
(0.001) (0.001)
familiarity×$\times$×count 0.088∗∗∗ 0.194∗∗∗
(0.014) (0.026)
status×$\times$×count 0.000 0.001∗∗
(0.000) (0.000)
familiarity×$\times$×status×$\times$×count -0.003∗∗∗
two-way fixed effects Y Y Y Y Y Y
Observations 11,510 16,877 16,615 11,359 11,359 11,359
N. of groups 2,817 5,360 5,309 2,804 2,804 2,804
$R^{2}$ 0.287 0.156 0.609 0.671 0.689 0.695
Residual Std. Error 0.743 0.497 0.987 1.141 1.156 1.161
F Statistic 1747.213∗∗∗ 2119.285∗∗∗ 5849.378∗∗∗ 2904.998∗∗∗ 2099.285∗∗∗ 1941.321∗∗∗
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Research persistence shows contrasting effects on productivity and impact. It is positively associated with productivity both before and after the paradigm shift - researchers who remained committed to their established topics or methods tended to publish more papers. Adaptive researchers experiencing a temporary decline in productivity could be due to several factors: they may have begun submitting to ICLR, a conference focused on deep learning, or needed time to regain their previous productivity levels while learning new methods and adjusting to the emerging paradigm (see Appendix C for more discussions). On the other hand, the interaction term between persistence and the post-2013 period reveals a significant negative relationship with impact, indicating that maintaining early research directions hindered the production of influential work after the deep learning boom. In the era following the advent of deep learning, given identical coauthor networks and prior performance, a 0.1 increase in similarity to prior research can explain a 1.77% marginal decrease in impact.
Collaborator familiarity, count, and status exhibit mixed effects on productivity and impact. Having more collaborators enhances productivity and impact. Collaborating with high-status partners appears to significantly enhance influence. However, this advantage comes with a tradeoff, as such collaborations tend to negatively affect productivity, potentially due to the time and resources required to maintain these high-status relationships and high-status authors’ high standard of publishing. Teams with older collaborations exhibit lower productivity and impact levels. Old teams are less successful in adapting to the new paradigm. We further examined the ”age” [^6] This age is the age of a team. In the paper we also have the age of a scientist, defined as the years since their first publication. of these scientific teams as shown in Figure 4, defined as the average number of ICML/NeurIPS papers the focal author had previously collaborated on with team members that year (right), and the proportion of new collaborations, defined as the percentage of team members who had never co-authored an ICML/NeurIPS paper with the focal author before (left). We tracked their collaborations back to the initial year of these conferences. We found that right after the shock, published teams tended to be more experienced — presumably those at the center of deep learning prior to the shock - and they continued to adapt through collaborating with newcomers in the years following 2013. This is evidenced by a decreasing average team age and an increasing share of new collaborators, both of which contributed to the continued success. Notably, these new collaborations tended to persist over time, as teams gradually became old again in subsequent years. These effects are robust within individual conferences, see Appendix B.
Figure 4: Teams adapt through fresh collaborations. Error bars represent 95% confidence intervals.
Another noteworthy finding is that prior impact has a significant suppressive effect on subsequent impact. Productivity shows a similar, although not statistically significant, trend after controlling the interaction effects between features of the collaboration network. In other words, a smaller prior impact is associated with a greater subsequent impact. This phenomenon is highly counterintuitive and suggests that during the turbulent period when deep learning emerged and became widely adopted, the influence of scientists was redistributed.
What drives this redistribution of influence? One possible mechanism is natural aging, which may explain why scientists’ early influence inversely predicts their future influence. In the regression analysis, we controlled for authors’ fixed effects to account for individual characteristics, such as age, motivation, and adaptivity when confronting the deep learning wave. Over the (brief) decade of post-deep-learning we tracked, it is possible that some scientists’ drive increased (e.g., those in their career growth phase), while others’ drive declined (e.g., those who were already well-established). In Appendix E, we track productivity and impact by academic age (the years since their first publication), controlling author fixed effects, across the cohort of scientists and find no strong evidence that impact or productivity decreases with age. In fact, while there is no meaningful effect on impact, productivity actually increases with age.
5.2.2 Who will be more adaptive?
Who is more likely to adapt? Adaptivity is pivotal to both impact and productivity in the post-deep learning era. If previously high-status scientists become less adaptive, they will lose status which could result in a redistribution of influence as observed in Section 5.2.1. This section explores this critical issue by examining the features of authors and their adaptiveness. We test it using the following regression:
$\textit{Research Persistence}_{i,t}=\;$ $\beta_{1}\times\textit{Prev. Impact}_{i,t}+\beta_{2}\times\textit{Prev. Productivity}_{i,t}$
$+\beta_{3}\times\textit{Collaborator Familiarity}_{i,t}+\beta_{4}\times\textit{Collaborator Count}_{i,t}$
$+\beta_{5}\times\textit{Collaborator Status}_{i,t}$
[…]
$+\alpha_{i}+\gamma_{t}+\varepsilon_{i,t}$
Table 3: Regression Results: Persistence as Dependent Variable
Dependent variable: persistence
(1) (2) (3) (4)
prev. productivity 4.040∗∗∗ 3.236∗∗∗
(0.150) (0.127)
prev. impact 0.019∗∗ 0.015∗∗
(0.006) (0.006)
collab. familiarity 1.828 1.995∗∗
(0.772) (0.738)
collab. status 0.008 0.008
(0.012) (0.011)
collab. count 0.618∗∗∗ 0.635∗∗∗
(0.107) (0.101)
familiarity_status -0.002 -0.002
(0.014) (0.014)
familiarity_count 0.833∗∗∗ 0.503
(0.213) (0.207)
count_status 0.003 -0.000
(0.003) (0.003)
familiarity_count_status -0.005 -0.011
(0.004) (0.004)
two-way fixed effects Y Y Y Y
Observations 11,510 11,319 11,359 11,170
N. of groups 2,817 2,787 2,804 2,772
$R^{2}$ 0.234 0.001 0.193 0.305
Residual Std. Error 5.825 0.444 5.279 6.636
F Statistic 2,652.971∗∗∗ 11.620∗∗∗ 290.508∗∗∗ 408.376∗∗∗
(df=2,836; 8,674) (df=2,806; 8,513) (df=2,829; 8,530) (df=2,799; 8,371)
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Our analysis reveals that a scientist’s past influence and productivity play a significant role in shaping the trajectory of their research evolution. Prolific scholars tend to be less inclined to shift their research paradigms. Older and larger teams are less likely to adapt. These findings suggest that in a dynamic scientific environment, researchers who adopt high-yield strategies — particularly those characterized by visible and immediate signals such as productivity or collaboration with established, large teams — become increasingly reluctant to abandon them, while such persistence harms their future performance.
Figure 5: The relationship between persistence, status, and status change using linear fits. Error bands represent 95% confidence intervals.
We present the relationship between status and persistence more intuitively. To measure change, we calculated the ratios of productivity and impact relative to the previous year. Figure 5 illustrates how impact and productivity evolve with increasing persistence (top row) and how persistence varies with previous productivity and impact (bottom row). A clear trend emerges: as persistence increases, impact tends to decrease while productivity increases. Notably, the highest increase in impact does not occur when the research trajectory is entirely altered (persistence = 0). Instead, it is observed within a modest range of persistence values, corresponding to a 25-50% overlap with prior work (top row). Additionally, as previous impact increases, persistence shows a weak upward trend, whereas previous productivity strongly correlates with increased persistence (bottom row).
5.3 Summary of Results
Together, the results (Figure 6) from Section 6 highlight the dual-edged nature of disruption in scientific progress. On the one hand, periods of disruption—such as the advent of deep learning—shake up the established order, challenging traditional hierarchies and creating opportunities for adaptive scientists to break from entrenched paradigms. This process reveals a tension between persistence and adaptability: while researchers who maintain continuity in their work may benefit from greater productivity, they often struggle to achieve high impact in rapidly shifting landscapes. On the other hand, those willing to adapt their research focus, especially during paradigm shifts, may be better positioned to produce influential contributions, despite the risks and costs associated with such changes.
Figure 6: Summary of key results.
Interestingly, our results suggest that the motivations and capacities for adaptation vary across different strata of the scientific community. Low-status scientists, constrained entrenched positions, may feel a heightened urgency to explore new directions during periods of disruption. This allows them to capitalize on emerging opportunities and challenge the existing power dynamics. This interplay of disruption, status, and adaptability underscores the complexity of scientific evolution, where both the persistence of existing paradigms and the emergence of novel directions coexist and compete.
6 Discussions
This study examines the dynamics of scientific progress during a major paradigm shift in the field of machine learning, focusing on how researchers’ strategies between persistence and adaptability could lead to future success. While persistence is positively associated with productivity, during the post-2013 deep learning era, it is negatively associated with impact. This finding suggests that rigid adherence to earlier research directions, especially during a major paradigm shift, hinders a scientist’s ability to produce influential work in rapidly evolving fields.
Notably, the highest impact is observed among researchers who achieve a balance between persistence and adaptation. Those who moderately alter their research direction, integrating elements of the new paradigm while retaining some continuity, are better positioned to produce impactful contributions. This finding underscores the importance of strategic adaptation, where researchers selectively pivot their focus to align with emerging trends without completely abandoning their prior expertise. Compared to empirical evidence that adaptive researchers are often penalized in the scientific system [35, 45, 55, 27], we believe our slightly different conclusion is drawn from a special paradigm-shifting event, where the relationship between adaptation and impact is more contextual and dynamic. As the community changes within the paradigm shift, so does the criteria of research evaluation.
The interplay between author status and adaptability further highlights the redistributive effects of disruption. Low-status scientists, often constrained by limited resources and recognition, are more likely to adapt during periods of disruption, seizing opportunities to reposition themselves within the scientific hierarchy. Conversely, high-status scientists and large, established teams, while better resourced, face stronger incentives to maintain their existing paradigms, which can hinder their ability to adapt quickly. This dynamic creates a redistribution of impact, where periods of disruption provide opportunities for some while reinforcing barriers for others. Such situations challenge traditional hierarchies and create opportunities for innovation by forcing researchers to reevaluate their approaches, ultimately benefiting those who can adapt quickly through fresh collaborations. As a result, disruption acts as both a destabilizing force and a catalyst for innovation, driving the advancement of knowledge by pushing scientists to confront the limits of their existing approaches.
Beyond findings at the micro-level, we also observe how ICML and NeurIPS conferences adapted to the deep learning paradigm, increasingly aligning with the cutting-edge topics emphasized at ICLR. Trends in topic importance demonstrate a convergence among these top conferences, possibly achieved by active adaptation through peer review and selection processes, as well as passive adaptation through cross-conference submissions. This adaptation highlights how leading scientific venues respond to disruptive changes to maintain their relevance and influence. The alignment of ICML and NeurIPS with ICLR reflects broader institutional flexibility to incorporate new paradigms, a critical factor for sustaining their roles as flagship venues.
Our findings have important implications for individual researchers, institutions, funding agencies, and future research evaluation exercises. For researchers, the ability to balance persistence and adaptation is critical for maintaining relevance and achieving impact during periods of disruption. Institutions and funding agencies should recognize the value of both research strategies, providing support for researchers to explore new directions while maintaining their foundational expertise. Additionally, our findings raise a question about the nature of research evaluation: research articles often lack visible signals for quality, and researchers frequently rely on heuristics such as citation counts or authors’ status [7, 8] when they cite. As the trendiness of a topic is another heuristic that could boost citations, its relationship with the quality of work is often highly blurred during a major shift in the research paradigm. Effectively evaluating the quality of publications and evaluating the appropriate emphasis on productivity versus impact in evolving academic landscapes remain a critical challenge for all scientific communities.
Despite our timely contribution to understanding the persistence of scientific careers, our study has a few limitations that readers should note and that future work should address. This research only considered published papers due to the limited availability of submission data, particularly in the earlier years of these conferences. Future work could offer a more comprehensive view by examining submitted papers — how they adapt and how that adaptation influences success. Notably, adaptive mechanisms may also occur during the review process, meaning that analyzing published papers captures only part of the story. The development of open peer review infrastructures may provide access to such review data in the future (existing review data is only available in most recent years), enabling more thorough research on this topic. It is also worth studying how selected researchers published in other fields and in venues that were slow to accept the new paradigm.
Acknowledge: We thank Hongyuan Xia, Yufeng Xia, Nadav Kunievsky, James A. Evans, participants at the asis&t MET-STI workshop, and members of the UChicago Knowledge Lab for valuable discussions. This research was supported in part by the University of Chicago Research Computing Center. All errors are our own.
Appendix
Appendix A Analysis of the dataset of 580 thousand general machine learning papers
In this study, we focused on NeurIPS and ICML as a relatively controlled environment to explore whether researchers who had consistently published in top-tier conferences could maintain their success without adapting to new trends. To compare NeurIPS/ICML with other research in the same field, we analyzed 579,973 papers from the Microsoft Academic Graph (MAG) dataset categorized under ”machine learning” or ”deep learning,” spanning the same 20-year period that marked the formation and growth of deep learning. Using the SPECTER2 model [10, 47], we generated yearly embeddings for these papers to examine the distinctions between deep learning (approximated by ICLR (2013–2022)), ICML/NeurIPS, and general machine learning papers (2003–2022). Unlike word-level models like Word2Vec [41], which fail to capture the overall semantics of abstracts, or general-purpose sentence-level models such as Sentence-BERT [43], which lack domain-specific adaptation, or even scientific text-focused models like SciBERT [12], which do not incorporate citation information, SPECTER2 offers a distinct advantage. Pre-trained on large-scale scientific literature, SPECTER2 uses an encoder-only transformer architecture with self-attention to capture contextual relationships, enhanced by citation-based contrastive learning to encode domain-specific semantic links such as thematic similarity. These features make SPECTER2 exceptionally well-suited for generating nuanced sentence-level embeddings in scientific contexts, outperforming most existing representation methods [47]
To prepare the data, we tokenized the full titles and abstracts of papers with a maximum length of 512 tokens—sufficient to cover the vast majority of content—and processed the data in batches of 32. The [CLS] token embedding from the final transformer layer was used as a pooled representation, effectively capturing the semantic meaning at the abstract level.
Figure 7: The relationship between general machine learning research, ICLR, and ICML+NeurIPS.
Figure 8: Topical dynamics are different in ICML/NeurIPS and general machine learning.
We began by identifying the central embedding for ICLR, ICML/NeurIPS, and general machine learning papers, then collected the 500 nearest papers around each center to represent each cluster. Pairwise cosine similarities were calculated 5,000 times between clusters (one in 500 neighbors to one in another 500) to approximate the distribution of distances and their yearly trends. To represent the most papers within a group we use the median of each dimension of the embedding.
The results in Figure 7 reveal several notable insights. Before the rise of deep learning, NeurIPS and ICML papers already showed limited similarity to general machine learning papers (left). Following the establishment of ICLR, as anticipated, NeurIPS and ICML became increasingly aligned with ICLR, while other machine learning papers drifted further away, despite all being part of the broader machine learning field.
What makes ICML and NeurIPS special? We tracked the proportion of six classic traditional machine learning topics appearing in the abstracts of each group. As shown in Figure 8, we found that trends in ICML and NeurIPS seem to predict the trends of these topics in the general corpus. In other words, the decline of traditional machine learning topics occurred in top conferences before it was reflected in other texts. This is also why, in the main text, we consider these ICML/NeurIPS conference papers of a cohort of scientists — if the lower bar is set, the productivity of scientists almost always increases, and other venues are just not as sensitive as ICML/NeurIPS to the paradigm shift. By controlling a relatively stable and selective environment that well represents the frontier of deep learning research, we can examine the impact of adaptation on scientists’ productivity and influence.
Appendix B Teams adapt through fresh collaborations within the individual conference
Figure 9: Teams adapt through fresh collaborations within ICML.
Figure 10: Teams adapt through fresh collaborations within NeurIPS.
As shown in Figures 9 and 10 the patterns reported in the main paper (Figure 4) are robust within individual conferences.
Appendix C The impact of ICLR
The establishment of ICLR may influence the cohort of scientists publishing in ICML and NeurIPS in two ways.
Figure 11: The ratio of new authors.
The first possible channel is that the establishment of ICLR in 2013 (year 1) brought more attention to ICML and NeurIPS conferences, given their shared themes, leading to heightened competition and the replacement of some scientists. However, we find that the ratio of new authors (defined as those who did not appear in any prior year) in ICML and NeurIPS in the first few years after year 1 was not significantly higher than in previous and following years (see Figure 11).
The second possible channel is that some scientists shifted their focus to publishing at ICLR so they disappeared from ICML/NeurIPS. However, only 1.21% (65) of the focal group of scientists submitted at least one paper to ICLR 2013, yet they accounted for 65.7% of submissions and 77.3% of accepted papers at the conference. During the first five years of ICLR (2013–2017) - a period marked by heightened dynamism as scientists formed new collaborations to adapt (see Figures 4, 9, and 10) - papers published in ICLR accounted for only 6.8% of this cohort’s total publications. Although ICLR later grew in prominence, it is important to recognize that such shifts in impact take time to materialize.
Adaptive researchers indeed experience a temporary decline in productivity for complex reasons, such as a shift in attention (adaptive authors may send papers to more adaptive venues, like ICLR). However, the primary cause, as we analyzed above, should be likely the time required to regain their prior productivity while learning new methods and adjusting to the emerging paradigm. If we included ICLR publications in our analysis, those who adapted would likely benefit even more, since ICLR was essentially created for them. Our major conclusion (”rigidity penalty”) would be even stronger. But what’s more interesting is that even within a publication environment these scientists were already familiar with - one that actively tries to capture the cutting edge - there are still penalties for failing to adapt.
Appendix D Relationships between previous status, collaboration networks, and age
Figure 12: Relationship between previous status, collaboration networks, and age.
To clarify the relationships between the variables, Figure 12 presents the pairwise correlations among scientists’ research persistence, prior impact, prior productivity, age, and features of their collaboration networks (including collaborator count, collaborator status, and collaborator familiarity). Overall, the correlations are relatively weak. There is a moderate correlation between research persistence, previous productivity, and current collaborators, but we find no evidence of multicollinearity.
Appendix E Regressions of age and productivity/impact
To what extent are the declines in productivity and impact simply natural consequences of aging? In this section, we argue that, within our study population, age is not a primary factor influencing research output.
Figure 13: Relationship between age and productivity/impact.
Impact has a negative relation with age but the association is very weak (coefficient $=-0.0002$ (in a 0-100 scale), p = 0.048). Productivity actually has a strong positive relation with age (coefficient = 0.007, p = 0.000).
Appendix F TF-IDF token extraction from abstract
Raw Abstract 10 Extracted Tokens and Their TF-IDF Scores When we learn a new motor skill, we have to contend with both the variability inherent in our sensors and the task. The sensory uncertainty can be reduced by using information about the distribution of previously experienced tasks. Here we impose a distribution on a novel sensorimotor task and manipulate the variability of the sensory feedback. We show that subjects internally represent both the distribution of the task as well as their sensory uncertainty. Moreover, they combine these two sources of information in a way that is qualitatively predicted by optimal Bayesian processing. We further analyze if the subjects can represent multimodal distributions such as mixtures of Gaussians. The results show that the CNS employs probabilistic models during sensorimotor learning even when the priors are multimodal. {sensorimotor: 0.467, sensory: 0.346, variability: 0.231, multimodal: 0.222, distribution: 0.204, cns: 0.19, subject: 0.188, represent: 0.172, task: 0.169, internally: 0.164}